A rock is dropped from a sea cliff and the sound of it striking the ocean is heard 3.6s later.
If the speed of sound is 340 m/s, how high is the cliff?
Answer #1

let the height of cliff is x

time taken for rock to fall , t = sqrt(2h/g)

t = sqrt(2x/g)

time taken for the sound to come up , t2 = distance/sped

t2 = x/340

Now , for the total time

sqrt(2x/9.8) + x/340 = 3.6

solving for x

x = 57.7 m

the height of the cliff is 57.7 m